A life saver has a 3cm diameter, a 1cm diamter hole and has 1/2 cm height. Find the volume and total surface area of the life saver. (Just like the candy lifesaver) That is what the pic looks like.
Thankyou Shelly

Answer by ankor@dixie-net.com(12692)

A life saver has a 3 cm diameter, a 1 cm diameter hole and has 1/2 cm height. Find the volume and total surface area of the life saver. (Just like the candy lifesaver) That is what the pic looks like.
:
The approach to this problem:
:
Use the volume (V1) of a cylinder (V=pi*r^2*h) to find the volume if it did not have a hole
Then find the volume (V2)of the material removed to form the hole, also a cylinder. Subtract V2 from V1
:
r = 1.5; h = .5
V1 = pi*1.5^2*.5
V1 = 3.5343 cu/cm
:
r = .5; h = .5
V2 = pi*.5^2*.5
V2 = .3927 cu/cm
:
3.5343 - .3927 = 3.1416 cu/cm is the volume of the life saver
:
:
Surface area approach:
:
Find the surface area of a cylinder, ignoring the hole.
Find area of the 2 circles that form the hole. Subtract from the total surface area.
Find the area of the sides of the cylinder forming the hole. Add to above area
:
Surface area, ignoring the hole
r = 1.5, d = 3, h = .5
SA = 2(pi*1.5^2) + (pi*3*.5)
SA = 14.1371 + 4.7124
SA = 18.8495 sq/cm
:
Area of the 2 circles that form the hole
A = (pi*.5^2) * 2
A = 1.5708 sq/cm
:
18.8495 - 1.5708 = 17.2787 sq/cm, surface area not including the inside of the hole
:
Find the area of the sides inside the hole: d = 1; h = .5, add to the SA
A = pi*1*.5
A = 1.5708
:
Find total surface area:
17.2787 + 1.5708 = 18.8495 sq/cm, total surface area.
:
:
Did this makes sense, by the way what happens to the candy removed from the life saver to form the hole?
I know that with donuts, you can buy donut holes at the bakery here.
:
I'm sure, Shelly, these comments, interest you tremendously, hope this helps anyway.